a gaseous reaction, A2(g) =B(g) + 1/2C(g)      shows increase in pressure from 100 mm to 120 mm in 5 min.The rate of disaapearnce of A2.?

Grade:12

6 Answers

Naveen Kumar
9 years ago
Given the reaction is
..................................A2(g) …..............=......B(g)...... +....... 1/2C(g)
..............................100mm.......................0........................0 ….(initially)
................................100-p.....................p....................p/2....(after 5 minutes)
total pressure=(100-p)+p+p/2=120mm(given)
p/2=20
p=40mm
so rate of disappearence of A2=40mm/5min
and if we know the temperature and volume of the mixture, we can alsom calculate the amount(mole) of A2 disappeared .
.
Shriya Mehrotra
9 years ago
sir but the ans is 8 mm/min.
Shriya Mehrotra
9 years ago
yes sir got it..thanks..
Alice
6 years ago
Why is the rate.of dissapearnce of A is 40/5 and not 100-40/5 ???? Since we wrote A after 5 minutes to be 100-p
Sunita mishra
5 years ago
Because rate of rxn ie change in pressure by time taken.and total pressure provided by evey gas is 120 atm.which is generated by produt and reactant both.100-40/5 is the pressure created by A only..and rate is equal for reactant and products both.after putting 100-40/5 we get 12 which is not same for all.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Given the reaction is
A2(g) = B(g) + 1/2C(g)
100mm 0 0 (initially)
100-p p p/2 (after 5 minutes)
total pressure = (100-p) + p + p/2 =120mm(given)
p/2 = 20
p=40mm
so rate of disappearence of A2 = 40mm/5min
and if we know the temperature and volume of the mixture, we can alsom calculate the amount(mole) of A2 disappeared .

Thanks and Regards